OpenStudy (anonymous):

can anybody solve this : if 1,a1,a2....a(n-1) are nth roots of unity,then the value of (1-a1) (1-a2)(1-a3)...(1-a(n-)) is equal to ?

OpenStudy (anonymous):

Hello, @Reeti1 :)

OpenStudy (anonymous):

(:

OpenStudy (anonymous):

HI

OpenStudy (anonymous):

What's up?

OpenStudy (anonymous):

NM ! YOU SAY ? your asl ?

OpenStudy (anonymous):

14 I'm a boy England Now, let's get started on this question :)

OpenStudy (anonymous):

ahh this is 12th grade mathematics dude ;)

OpenStudy (anonymous):

Well, if you say so :3

OpenStudy (anonymous):

can you solve it ?

OpenStudy (anonymous):

Of course I can :P

OpenStudy (anonymous):

plaese move then

OpenStudy (anonymous):

Moving...

OpenStudy (anonymous):

move on*

OpenStudy (anonymous):

Okay, all the nth roots of unity are roots of this polynomial... \[\huge x^n - 1 = 0\]

OpenStudy (anonymous):

OKAY

OpenStudy (anonymous):

Consequently, this polynomial is equal to... \[\LARGE (x-1)(x-a_1)(x-a_2)...(x-a_{n-1})=0\]

OpenStudy (anonymous):

But also \[\LARGE x^n - 1 = (x-1)(x^{n-1}+x^{n-2}+...x+1)\]

OpenStudy (anonymous):

Sorry \[\Large (x-a_1)(x-a_2)...(x-a_{n-1})=x^{n-1}+x^{n-2}+...x+1\]

OpenStudy (anonymous):

Is everything clear? @Reeti1

OpenStudy (anonymous):

the choices are : a. 1 b. 1/2 c. n

OpenStudy (anonymous):

Yeah yeah... but everything is clear to you so far? How we arrived at this... \[\Large (x-a_1)(x-a_2)...(x-a_{n-1})=x^{n-1}+x^{n-2}+...x+1\]

OpenStudy (anonymous):

answer is n i guess

OpenStudy (anonymous):

YES! See? Just replace x with 1 \[\Large (1-a_1)(1-a_2)...(1-a_{n-1})=1^{n-1}+1^{n-2}+...1+1\]

OpenStudy (anonymous):

And you get n 1's that you'll add to get... n haha did I solve it or what? ^.^

OpenStudy (anonymous):

you did it ! but wow how ? you're just 14

OpenStudy (anonymous):

I'm actually well over 200? ^.^ @dmezzullo already said on chat... age doesn't matter XD

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

ok one more

OpenStudy (anonymous):

Let's go :)

OpenStudy (anonymous):

cube root of i is equal to ?

OpenStudy (anonymous):

cube root of i? :( \[\huge \sqrt[3]i\]

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

is that even possible? :/

OpenStudy (anonymous):

haha of course it is :D

OpenStudy (anonymous):

First, put i in polar form ^.^

OpenStudy (anonymous):

okay done